# Human-in-the-Loop Causal Discovery under Latent Confounding using Ancestral GFlowNets

### v2

@@ Abstract
Structure learning is the crux of causal inference. Notably, causal discovery (CD) algorithms are brittle when data is scarce, possibly inferring imprecise causal relations that contradict expert knowledge — especially when considering latent confounders. To aggravate the issue, most CD methods do not provide uncertainty estimates, making it hard for users to interpret results and improve the inference process. Surprisingly, while CD is a human-centered affair, no works have focused on building methods that both 1) output uncertainty estimates that can be verified by experts and 2) interact with those experts to iteratively refine CD. To solve these issues, we start by proposing to sample (causal) ancestral graphs proportionally to a belief distribution based on a score function, such as the Bayesian information criterion (BIC), using generative flow networks. Then, we leverage the diversity in candidate graphs and introduce an optimal experimental design to iteratively probe the expert about the relations among variables, effectively reducing the uncertainty of our belief over ancestral graphs. Finally, we update our samples to incorporate human feedback via importance sampling. Importantly, our method does not require causal sufficiency (i.e., unobserved confounders may exist). Experiments with synthetic observational data show that our method can accurately sample from distributions over ancestral graphs and that we can greatly improve inference quality with human aid.
@@ 1 Introduction
Drawing conclusions about cause-and-effect relationships presents a fundamental challenge in various scientific fields and significantly impacts decision-making across diverse domains Pearl (2000). The importance of having structural knowledge, often encoded as a causal diagram, for conducting causal inferences is widely recognized, a concept made prominent by Cartwright (1989)’s dictum: “no causes in, no causes out”. When there is no objective knowledge to fully specify a causal diagram, causal discovery (CD) tools are instrumental in partially uncovering causal relationships among variables from, for example, observational data. Formally, let $\mathbf{V}=\{V_{1},V_{2},\dots,V_{n}\}$ be a set of $n$ observed variables and $\mathcal{D}$ be a dataset containing $|\mathcal{D}|=m$ samples for each $V_{i}\in\mathbf{V}$. A CD algorithm takes $\mathcal{D}$ as input and typically returns a single graph $\mathcal{G}=(\mathbf{V},\mathbf{E})$ with well-defined causal semantics, in which each node in $\mathbf{V}$ represents a variable $V_{i}\in\mathbf{V}$ and each edge in $\mathbf{E}$ encodes the possible underlying (causal/confounding) mechanisms compatible with $\mathcal{D}$.
Figure 1: Human-in-the-loop probabilistic CD. We first train an AGFN to fit a data-informed belief over AGs. Then, we iteratively refine it by 1) questioning (Q) experts on the relation between a highly informative pair of variables and 2) updating the belief given the potentially noisy answers (A). The histograms on top of the edges show marginals over edge types (green denotes ground truth). Notably, our belief increasingly concentrates on the true AG, $1\rightarrow 2\leftrightarrow 3$.
This work focuses on recovering the structure of the underlying causal diagram when unobserved confounders may be at play. We propose to address this task by not only leveraging observational data but also by accounting for potentially noisy pieces of expert knowledge, otherwise unavailable as data. Throughout this work, we consider ancestral graphs (AGs) as surrogates for causal diagrams. AGs are particularly convenient since they encode latent confounding without explicitly invoking unobserved variables. Moreover, AGs capture all conditional independencies and ancestral relations among observed variables $\mathbf{V}$, as entailed by a causal diagram (Richardson and Spirtes, 2002).
In the realm of CD from solely observational data, algorithms aim to construct a compact representation of the joint observational distribution $P(\mathbf{V})$, which implies a factorization as a product of conditional probabilities. Notably, multiple models may entail the same conditional independencies; in such cases, they are denoted as Markov-equivalent. As a result, these algorithms can only reconstruct the class of Markov-equivalent models (AGs), denoted as the Markov Equivalence Class (MEC) and typically represented by a Partial Ancestral Graph (PAG). Importantly, CD beyond the MEC by leveraging domain knowledge presents a critical challenge. Notably, there is no proper characterization of an equivalence class that accounts for knowledge stemming from both humans and data (Wang et al., 2022).
There is a variety of algorithms for CD from observational data, primarily categorized into constraint- and score-based methods. The former uses (in)dependence constraints derived via conditional independence tests to directly construct a PAG representing the MEC. The latter uses a goodness-of-fit score to navigate the space of AGs, selecting an optimum as a representative for the MEC.
Nonetheless, methods within both paradigms suffer from unreliability when data is scarce. Specifically, for the majority of the CD algorithms, formal assurances that the inferred MEC accurately represents the true causal model heavily rely on the so-called faithfulness assumption, which posits that all conditional independencies satisfied by $P(\mathbf{V})$ are entailed by the true causal model (Zhang and Spirtes, 2016). However, this presents a critical challenge in real-world scenarios, as violations of the faithfulness assumption become more prominent when relying on $P(\mathbf{V})$ estimated from limited data Uhler et al. (2012); Andersen (2013); Marx et al. (2021). For constraint-based methods, hypothesis tests may lack the statistical power to detect conditional independencies accurately. These errors may propagate and trigger a chain reaction of erroneous orientations Zhang and Spirtes (2008); Zhalama et al. (2017b); Ng et al. (2021). For score-based methods, although score functions directly measure goodness-of-fit on observational data, small sample sizes can significantly skew the estimates for the population parameters. Consequently, structures deemed score-optimal may not necessarily align with the ground-truth MEC  Ogarrio et al. (2016). A major concern is that the overwhelming majority of CD algorithms produce a single representation of the MEC as output, without quantifying the uncertainty that arises during the learning process Claassen and Heskes (2012); Jabbari et al. (2017). This poses a significant challenge for experts, as it hinders their ability to validate the algorithm’s outcome or gain insights into potential venues for improving inference quality.
To alleviate the lack of uncertainty quantification in CD, we propose sampling AGs from a distribution defined using a score function, which places best-scoring AGs around the mode by design. This effectively provides end-users with samples that reflect the epistemic uncertainty inherent in CD, thus allowing their propagation through downstream causal inferences. In particular, we sample from our belief using Generative Flow Networks (GFlowNets; Bengio et al., 2021a, b), which are generative models known for sampling diverse modes while avoiding the mixing time problem of MCMC methods, and not requiring handcrafted proposals nor accept-reject steps (Bengio et al., 2021a).
Acknowledging the low-data regime as CD’s primary challenge, we also propose actively integrating human feedback in the inferential process. This involves modeling user knowledge on the existence and nature (confounding/ancestral) of the relations and using it to weigh our beliefs over AGs. During our interactions with experts, we probe them about the relation of the pair of variables that optimizes a utility/acquisition function, for which we propose the negative expected cross-entropy between our prior and updated beliefs. Unlike prior strategies, our acquisition avoids the need to estimate the normalizing constant and predictive distribution of our updated belief, as needed for information gain and mutual information, respectively. Notably, we use importance sampling (Marshall, 1954; Geweke, 1989) to update our initial belief with the human feedback, which avoids retraining GFlowNets after each human interaction.
While incorporating expert knowledge into CD has been a long-standing goal (Meek, 1995a; Chen et al., 2016; Li and Beek, 2018; Wang et al., 2022), existing works either rely on strong assumptions (e.g., causal sufficiency) or assume the knowledge is noiseless, aligned with the ground truth  (Andrews, 2020; Wang et al., 2022). Importantly, our work introduces the first iterative CD framework for AGs involving a human in the loop and accommodating potentially noisy feedback, as depicted in Figure 1.
To validate our approach, we conduct experiments using the BIC score, for linear Gaussian causal models. Specifically, we assess: i) our ability to sample from score-based beliefs over AGs, ii) how our samples compare to samples from bootstrapped versions of state-of-the-art (SOTA) methods, and iii) the efficacy of our active knowledge elicitation framework using simulated humans. We observe that our method, Ancestral GFlowNet (AGFN), i) accurately samples from our beliefs over AGs; ii) consistently includes AGs with low structural error among its top-scored samples; and iii) is able to greatly improve performance metrics (i.e., SHD and BIC) when incorporating human in the loop.
In summary, the contributions of our work are:
@@
    We leverage GFlowNets to introduce AGFN, the first CD algorithm for scenarios under latent confounding that employs fully-probabilistic inference on AGs;
@@
    We show AGFN accurately learns distributions over AGs, effectively capturing epistemic uncertainty.
@@
    We propose an experimental design to query potentially noisy expert insights on relationships among pairs of variables that lead to optimal uncertainty reduction.
@@
    We show how to incorporate expert feedback into AGFN without retraining them from scratch.
@@ 2 Background
This section introduces the relevant notation and concepts. We use uppercase letters $V$ to represent a random variable or node in a graph, and boldface uppercase letters $\mathbf{V}$ to represent matrices or sets of random variables or nodes.
Ancestral graphs. Under the assumption of no selection bias, an ancestral graph (AG) $\mathcal{G}$ over $\mathbf{V}$ is a directed graph comprising directed ($\rightarrow$) and bidirected ($\leftrightarrow$) edges Richardson and Spirtes (2002); Zhang (2007). In any directed graph, if a sequence of directed edges, $V_{i}\rightarrow\cdots\rightarrow V_{j}$, connects two nodes $V_{i}$ and $V_{j}$, we refer to this sequence as a directed path. In this case, we also say that $V_{i}$ is an ancestor of $V_{j}$ and denote this relation as $V_{i}\in An(V_{j})$. By definition, any AG $\mathcal{G}$ must further satisfy the following:
@@
    there is no almost directed cycle, i.e., if $V_{i}\leftrightarrow V_{j}$ is in $\mathcal{G}$, then $V_{j}\not\in An(V_{i})$ and $V_{i}\not\in An(V_{j})$.
As a probabilistic model, the nodes in an AG represent random variables, directed edges represent ancestral (causal) relationships, and bidirected edges represent associations solely due to latent confounding. For a complete characterization of AGs, refer to Richardson and Spirtes (2002).
Data generating model. We assume that the data-generating model corresponds to a linear Gaussian structural causal model (SCM) (Pearl, 2000) defined by a 4-tuple $\mathcal{M}=\langle\mathbf{V},\mathbf{U},\mathcal{F},P(\mathbf{U})\rangle$, in which $\mathbf{V}=\{V_{1},V_{2},\dots,V_{n}\}$ is a set of $n$ observed random variables and $\mathbf{U}=\{U_{1},U_{2},\dots,U_{n}\}$ is the set of unobserved random variables. Further, let $Pa_{i}\subseteq\mathbf{V}\setminus\{V_{i}\}$ be the set of observed causes (parents) of $V_{i}$, and $U_{i}$ be the set of unobserved causes of $V_{i}$. Then, each structural equation $f_{i}\in\mathcal{F}$ is defined as:
@@
with $P(\mathbf{U})$ being a multivariate Gaussian distribution with zero mean and a not necessarily identity covariance matrix $\boldsymbol{\Omega}=(\omega_{ij})_{1\leq i,j\leq n}$ — the error terms $\{U_{i}\}$ are not necessarily mutually independent, implying that the system can be semi-Markovian (i.e., latent confounding may be present).
Consider a lower triangular matrix of structure coefficients $\mathbf{B}=(\beta_{ij})_{1\leq i,j\leq n}$ such that $(\mathbf{I}-\mathbf{B})$ is invertible, and $\beta_{ij}\neq 0$ only if $V_{j}\in Pa_{i}$. Then, the set of structural equations is given in matrix form by
@@
is represented by an AG in which, for every $i\neq j$, there is a directed edge $V_{j}\rightarrow V_{i}$ if $\beta_{ij}\neq 0$ and a bidirected edge $V_{j}\leftrightarrow V_{i}$ if $\omega_{ij}\neq 0$ (Richardson and Spirtes, 2002).
@@
Figure 2: Illustration of the generative process of AGs $\{\mathcal{G}_{5},\mathcal{G}_{6},\mathcal{G}_{7}\}$ using GFlowNets. Starting with an empty graph $\mathcal{G}_{0}$, we add edges between variables $\{X_{1},X_{2},X_{3}\}$ according to the action-policy $\pi_{F}$. Solid edges trace trajectories leading to sampled graphs. Dashed lines represent non-realized transitions to terminal state $\square$.
GFlowNets. Generative Flow Networks (GFlowNet; Bengio et al., 2021a, b) are generative models designed to sample from a finite domain $\mathcal{X}$ proportionally to some reward function $R:\mathcal{X}\rightarrow\mathbb{R}_{+}$, which may be parametrized using neural networks. In this work, we define $R$ as a strictly decreasing transformation of the BIC (more details in section 3). GFlowNets also assume there is a compositional nature to the elements $x\in\mathcal{X}$, meaning that they can be built by iteratively acting to modify a base object (i.e., an initial state). For instance, graphs can be built by adding edges to a node skeleton (Deleu et al., 2022) or molecules by adding atoms to an initial structure (Bengio et al., 2021a).
The generative process follows a trajectory of states $s\in\mathcal{S}$ guided by a transition probability $\pi_{F}:\mathcal{S}^{2}\rightarrow[0,1]$. In turn, $\pi_{F}$ is proportional to a foward flow function $F_{\theta}:\mathcal{S}^{2}\rightarrow\mathbb{R}_{+}$, which is parameterized by a neural network $\theta$. Let $\text{Pa}(s^{\prime})$ ($\text{Ch}(s^{\prime})$) be the set of all states which can transition into (directly reached from) $s^{\prime}$. Then, $\pi_{F}$ is defined as
@@
The support $\mathcal{X}$ of $R$ is contained within $\mathcal{S}$. There are also two special states in $\mathcal{S}$: an initial state $s_{0}$ and a terminal state $s_{f}$. We start with the initial state $s_{0}$ and transform it to a new valid state $s$ with probability $\pi_{F}(s|s_{0};\theta)$. We keep iterating this procedure until reaching $s_{f}$. States $s$ valid as final samples ($s\in\mathcal{X}$) are known as terminating states and have a positive probability for the transition $s\rightarrow s_{f}$. Figure 2 illustrates this process with $\mathcal{X}$ being the space of AGs. Crucially, the same parameterization $\theta$ is used for all transition probabilities $\pi_{F}(\cdot|s;\theta)$ given any departing state $s$, allowing for generalization to states never visited during training.
As the GFlowNet framework requires that no sequence of actions leads to a loop, we represent the space of possible action sequences by a pointed Directed Acyclic Graph (DAG) (Bengio et al., 2021b). The generation of any sample $x\in\mathcal{X}$ follows a trajectory $\tau=(s_{0},s_{1},\ldots,s_{T}=x,s_{f})\in\mathcal{S}^{T+2}$ for a $T\geq 0$. Different trajectories may lead to the same sample $x$. To ensure we sample proportionally to $R$, we search for a GFlowNet that satisfies the flow-matching condition, i.e., $\forall s^{\prime}\in\mathcal{S}$:
@@
Equation 5 implies the flow that enters $s^{\prime}$ equals the flow leaving $s^{\prime}$, except for some flow $R(s^{\prime})$ leaking from $s^{\prime}$ into $s_{f}$. We let $R(s)=0$ for $s\notin\mathcal{X}$. Eventually, it may be that all states $s$ are valid candidates, i.e., $\mathcal{S}=\mathcal{X}\cup\{s_{f}\}$. If so, each of eq. 5’s solutions satisfies a detailed-balance condition,
@@
for a parametrized backward flow $F_{B,\theta}\colon\mathcal{S}^{2}\rightarrow\mathbb{R}_{+}$ Deleu et al. (2022). In practice, we enforce eq. 6 by minimizing

### v3

# Expert-Aided Causal Discovery of Ancestral Graphs

@@ Abstract
Causal discovery (CD) algorithms are notably brittle when data is scarce, inferring unreliable causal relations that may contradict expert knowledge, especially when considering latent confounders. Furthermore, the lack of uncertainty quantification in most CD methods hinders users from diagnosing and refining results. To address these issues, we introduce Ancestral GFlowNets (AGFNs). AGFN samples ancestral graphs (AGs) proportionally to a score-based belief distribution representing our epistemic uncertainty over the causal relationships. Building upon this distribution, we propose an elicitation framework for expert-driven assessment. This framework comprises an optimal experimental design to probe the expert and a scheme to incorporate the obtained feedback into AGFN. Our experiments show that: i) AGFN is competitive against other methods that address latent confounding on both synthetic and real-world datasets; and ii) our design for incorporating feedback from a (simulated) human expert or a Large Language Model (LLM) improves inference quality.
@@ 1 INTRODUCTION
Causal discovery (CD) methods are widespread in science as tools to uncover complex cause-and-effect relationships. These algorithms typically leverage observational data to infer a graphical representation of the class of models that are equally likely to have generated the data, known as the Markov Equivalence Class (MEC). However, they are known to be unreliable when data are scarce, as statistical relationships drawn from such data may not align with the true causal model. This mismatch defines what we call a violation of the faithfulness assumption (Zhang and Spirtes,, 2016). For instance, CD algorithms based on conditional independence tests may fall victim to false independence relations inferred due to a lack of statistical power. These statistical errors may propagate and trigger a chain reaction of erroneous edge orientations (Zhang and Spirtes’,, 2008; Zhalama et al., 2017b, ; Ng et al.,, 2021). Meanwhile, algorithms that maximize goodness-of-fit scores may infer a structure that is optimal for the observed data, but does not align with the ground-truth MEC (Ogarrio et al.,, 2016).
In the presence of latent confounding, CD focuses on learning Ancestral Graphs (AGs), which conveniently encode causal structures while abstracting away unobserved confounders (Richardson and Spirtes,, 2002). Notably, learning AGs is particularly challenging as latent confounders both enlarge the space of models consistent with the data and weaken observable dependencies, thereby exacerbating model ambiguity and increasing the likelihood of faithfulness violations.
CD algorithms such as FCI (Zhang, 2008b, ), GFCI (Ogarrio et al.,, 2016), ACI (Magliacane et al.,, 2016), and DCD (Bhattacharya et al.,, 2021) address latent confounding but lack uncertainty quantification, which is crucial under limited samples and possible faithfulness violations. Some extensions incorporate expert knowledge (Ribeiro et al.,, 2024; Wang et al.,, 2023; Andrews,, 2020; Ankan and Textor,, 2025), yet they require deterministic, noise-free inputs and support only a narrow range of knowledge types. A complete and seamless integration of probabilistic or noisy expert information remains absent.
From a practitioner’s perspective, two questions are particularly relevant when interpreting CD results:
@@
    How much should results be trusted?
@@
    How can the quality of inference be improved?
Two key components to address these questions are: 1) integrating uncertainty quantification into CD to enhance trustworthiness and transparency; and 2) developing mechanisms to incorporate experts in the loop.
With this in mind, we propose a two-stage fully probabilistic CD framework called Ancestral GFlowNets (AGFNs). 1. First, an AGFN approximates an energy-based distribution over AGs based on a given score function such as the Bayesian Information Criterion (Schwarz,, 1978). 2. Then, it probes an uncertain expert (possibly an LLM) on the relationship between selected pairs of variables and uses their feedback for self-refinement. In doing so, AGFN can provide accurate responses even in the presence of unfaithful data, in contrast with standard CD algorithms.
While most probabilistic CD rely on Markov Chain Monte Carlo (MCMC) methods (e.g., Silva and Ghahramanir,, 2009; Silva,, 2013), AGFNs sample AGs using Generative Flow Networks (GFlowNets; Bengio et al., 2021a, ; Bengio et al., 2021b, ), which are generative models designed to sample diverse modes without the mixing time limitations of MCMC and without relying on handcrafted proposals nor accept-reject steps (Bengio et al., 2021a, ). Importantly, the score function is a hyperparameter of AGFN, allowing users to select whichever function is most suitable for their data without requiring changes in our method. This extends the applicability of AGFN beyond linear Gaussian data. Additionally, although GFlowNets have been previously used for CD (Deleu et al.,, 2022, 2023), AGFNs are the first method that simultaneously account for latent confoundedness of the data and allow for the seamless integration of expert knowledge into the discovery process. For a detailed discussion on other related works, we refer the reader to Appendix E.
To achieve this, we design an active elicitation strategy for AGFN for querying an expert about the existence and nature (whether ancestral or confounding) of an edge between the most informative pair of variables (Ryan et al.,, 2015). Subsequently, we update AGFN’s beliefs based on this (potentially noisy) feedback through a procedure that bypasses the need for model retraining. Figure 1 illustrates our iterative framework for expert-in-the-loop (EITL) probabilistic CD.
Figure 1: EITL probabilistic CD. We iteratively refine the trained AGFN by i) querying (Q) experts on the relationship of a highly informative pair of variables and ii) updating beliefs based on their (noisy) answers (A). Histograms show marginals over edge types (green denotes ground truth). Notably, our belief increasingly concentrates on the true AG, $1\rightarrow 2\leftrightarrow 3$.
Crucially, AGFN introduces a paradigm shift in causal inference by allowing the integration of uncertain knowledge in addition to the observational data. Incorporating background knowledge typically requires characterizing equivalence classes and designing specialized inference methods, but AGFN circumvents this by operating on distributions over AGs. Inference on each AG sample can be averaged, naturally propagating uncertainty.
To validate our approach, we conduct experiments using the BIC score for linear Gaussian causal models, assessing i) our ability to accurately sample from score-based beliefs over AGs, ii) how these samples compare to those obtained from bootstrapped versions of state-of-the-art (SOTA) CD methods, and iii) the effectiveness of our active knowledge elicitation framework with simulated humans and an LLM. We observe that our method, AGFN, i) accurately samples from our beliefs over AGs; ii) consistently includes AGs with low structural error among its top-scored samples; and iii) is able to greatly improve performance metrics (i.e., SHD and BIC) when incorporating experts in the loop.
In summary, our contributions are:
  1. 1.
    We propose AGFN, the first expert-in-the-loop probabilistic CD algorithm that supports uncertainty over the expert’s responses.
  2. 2.
    We develop an elicitation framework that enables AGFN to optimally interact with experts by iteratively selecting maximally informative questions. This allows inference refinement that combines observational information with expert knowledge.
@@
    We introduce a Bayesian algorithm to update AGFN’s distribution from expert responses, entirely bypassing model retraining.
@@
    We evaluate AGFN on diverse CD tasks, including the realistic benchmark Sachs dataset (Sachs et al.,, 2005). Results show that AGFN accurately and efficiently samples from its belief distribution and is competitive with SOTA CD in SHD and BIC. We also show that LLMs can be effectively used as surrogate experts for the EITL pipeline.
@@ 2 BACKGROUND
This section introduces the relevant notation and concepts. Uppercase letters ($V$) represent random variables or graph nodes. Bold uppercase letters ($\mathbf{V}$) represent matrices or sets of random variables or nodes.
2.1 Ancestral graphs
Assuming no selection bias, an ancestral graph (AG) $\mathcal{G}$ over $\mathbf{V}$ is a directed graph comprising directed ($\rightarrow$) and bidirected ($\leftrightarrow$) edges (Richardson and Spirtes,, 2002; Zhang,, 2007). In a directed graph, a sequence of directed edges $V_{i}\rightarrow\cdots\rightarrow V_{j}$ is called a directed path from $V_{i}$ to $V_{j}$. In this case, $V_{i}$ is an ancestor of $V_{j}$ and denote this relation as $V_{i}\in An(V_{j})$. By definition, any AG $\mathcal{G}$ must further satisfy the following:
@@ 2.1 Ancestral graphs
    there are no almost directed cycles, i.e., if $V_{i}\leftrightarrow V_{j}$ is in $\mathcal{G}$, then $V_{j}\not\in An(V_{i})$ and $V_{i}\not\in An(V_{j})$.
As a probabilistic model, an AG has nodes as random variables, directed edges as ancestral (causal) relations, and bidirected edges as associations solely due to latent confounding (Richardson and Spirtes,, 2002).
2.2 Linear Gaussian SCMs
A linear Gaussian SCM is defined by a 4-tuple $\mathcal{M}=\langle\mathbf{V},\mathbf{U},\mathcal{F},P(\mathbf{U})\rangle$, in which $\mathbf{V}=\{V_{1},\dots,V_{n}\}$ is a set of $n$ observed random variables and $\mathbf{U}=\{U_{1},\dots,U_{n}\}$ is the set of unobserved random variables. Further, let $Pa_{i}\subseteq\mathbf{V}\setminus\{V_{i}\}$ be the set of observed causes (parents) of $V_{i}$, and $U_{i}$ be the set of unobserved causes of $V_{i}$. Each structural equation $f_{i}\in\mathcal{F}$ is defined as:
@@ 2.2 Linear Gaussian SCMs
with $P(\mathbf{U})$ a multivariate Gaussian distribution with zero mean and a covariance matrix $\boldsymbol{\Omega}=(\omega_{ij})_{1\leq i,j\leq n}$ that is not necessarily the identity. The error terms $\{U_{i}\}$ are not necessarily mutually independent, implying that the system can include latent confounding.
Let $\mathbf{B}=(\beta_{ij})_{1\leq i,j\leq n}$ be a lower-triangular matrix of structural coefficients such that $(\mathbf{I}-\mathbf{B})$ is invertible, with $\beta{ij}\neq 0$ only if $V_{j}\in Pa_{i}$. The structural equations can then be expressed in matrix form as
@@
is represented by an AG with directed edges $V_{j}\rightarrow V_{i}$ if $\beta_{ij}\neq 0$ and bidirected edges $V_{j}\leftrightarrow V_{i}$ if $\omega_{ij}\neq 0$, for every $i\neq j$ (Richardson and Spirtes,, 2002). To validate AGFN in our synthetic experiments, we draw samples from a linear Gaussian SCM and compare the AGFN-learned AG distribution against the ground truth.
@@
Figure 2: Generative process of AGs $\{\mathcal{G}_{5},\mathcal{G}_{6},\mathcal{G}_{7}\}$ using GFlowNets. Starting from an empty graph $\mathcal{G}_{0}$, edges between variables $\{X_{1},X_{2},X_{3}\}$ are added following the action-policy $\pi_{F}$. Solid edges indicate trajectories leading to sampled graphs. Dashed lines denote non-realized transitions to the terminal state $\square$.
2.3 GFlowNets
GFlowNets are generative models designed to sample from a finite domain $\mathcal{X}$ proportionally to some reward function $R:\mathcal{X}\rightarrow\mathbb{R}_{+}$, which may be parametrized using neural networks. In this work, we define $R$ as a strictly decreasing transformation of the BIC (see more details in Section 3). GFlowNets also assume elements $x\in\mathcal{X}$ are compositional, built by iteratively modifying a base object (initial state). For instance, graphs can be generated by adding edges to an empty graph (Deleu et al.,, 2022), or molecules by adding atoms to an initial structure (Pandey et al.,, 2025).
The generative process follows a trajectory of states $s\in\mathcal{S}$ guided by a transition probability $\pi_{F}:\mathcal{S}^{2}\rightarrow[0,1]$. In turn, $\pi_{F}$ is proportional to a forward flow function $F_{\theta}:\mathcal{S}^{2}\rightarrow\mathbb{R}_{+}$, which is parametrized by a neural network $\theta$. Let $\text{Pa}(s^{\prime})$ ($\text{Ch}(s^{\prime})$) be the set of all states which can transition into (directly reached from) $s^{\prime}$. Then, $\pi_{F}$ is defined as
@@ 2.3 GFlowNets
The support $\mathcal{X}$ of $R$ is contained within $\mathcal{S}$. There are also two special states in $\mathcal{S}$: an initial state $s_{0}$ and a terminal state $s_{f}$. Starting from $s_{0}$, we iteratively transform it into a new valid state $s$ with probability $\pi_{F}(s\mid s_{0};\theta)$, repeating until reaching $s_{f}$. States $s$ that are valid as final samples ($s\in\mathcal{X}$) are known as terminating states and have a positive probability for the transition $s\rightarrow s_{f}$. Figure 2 illustrates this process with $\mathcal{X}$ being the space of AGs. The same parametrization $\theta$ is used for all transition probabilities $\pi_{F}(\cdot|s;\theta)$ given any departing state $s$, allowing for generalization to states not visited during training.
As GFlowNets do not allow action sequences to form loops, the space of action sequences is represented as a pointed Directed Acyclic Graph (DAG) (Bengio et al., 2021b, ). The generation of any sample $x\in\mathcal{X}$ follows a trajectory $\tau=(s_{0},\ldots,s_{T}=x,s_{f})\in\mathcal{S}^{T+2}$ with $T\geq 0$. Many trajectories may reach a same $x$. To sample proportionally to $R$, we require a GFlowNet satisfying the flow-matching condition, i.e., $\forall s^{\prime}\in\mathcal{S}$:
@@
Eq. (5) states that the flow into $s^{\prime}$ equals the flow out of $s^{\prime}$, except for some flow $R(s^{\prime})$ leaking from $s^{\prime}$ into $s_{f}$, with $R(s)=0$ for $s\notin\mathcal{X}$. When all states are valid samples, i.e., $\mathcal{S}=\mathcal{X}\cup\{s_{f}\}$, every solution of Eq.(5) satisfies a detailed-balance condition,
@@
for some backward flow $F_{B,\theta}\colon\mathcal{S}^{2}\rightarrow\mathbb{R}_{+}$ (Deleu et al.,, 2022). In practice, we enforce (6) by minimizing